OpenStudy (anonymous):

You graph a system of equations and the two lines overlap, or lie right on top of one another. Explain what this tells you about the: solution to the system the two equations the result if the system was solved algebraically

OpenStudy (anonymous):

Please @wolf1728 help me with this last one!

OpenStudy (anonymous):

Thanks wolf you came at the right time my mom is on my butt man i need to finish this last question quick

OpenStudy (wolf1728):

If the two lines lie one atop the other, the equations are said to be equivalent. (One equation can easily be transformed to the other by algebraic manipulation). Example of equivalent equations: x + y = 9 3x + 3y = 27 If two lines are parallel to each other, the equations are said to be inconsistent. Example 2x + 2y = 6 2x + 2y = 8 If the equations are equivalent or inconsistent, they cannot be solved. If equations are inconsistent, when solved algebraically they produce mathematical impossibilities. For example, solving the above 2 equations produces the result that 6 = 8. Solving equivalent equations: A) x + y = 9 B) 3x + 3y = 27 We'll multiply equation A by -3 A) -3x -3y = -27 and add that to B B) 3x + 3y = 27 which results in nothing.

OpenStudy (wolf1728):

Gee, I got a little carried away on that answer. :-)

OpenStudy (anonymous):

Lmao you did i'm not sure which part i should take for my answer

OpenStudy (wolf1728):

If the graph of both lines lie one atop the other, there will be no solution. The equations are said to be equivalent. If solved algebraically, the result is nothing. (Not too sure about that last sentence)

OpenStudy (wolf1728):

From above: Solving equivalent equations: A) x + y = 9 B) 3x + 3y = 27 We'll multiply equation A by -3 A) -3x -3y = -27 and add that to B B) 3x + 3y = 27 which results in nothing.

OpenStudy (anonymous):

Take that too?

OpenStudy (wolf1728):

That is a summary of the longer paragraph I typed. Yes, include the equivalent equation part.

OpenStudy (anonymous):

Okay so for this question I put If the graph of both lines lie one on top of the other, there will be no solution.The equations are said to be equivalent. If solved algebraically the result will be nothing. Solving equivalent equations: A) x + y = 9 B) 3x + 3y = 27 We'll multiply equation A by -3 A) -3x -3y = -27 and add that to B B) 3x + 3y = 27 which results in nothing.

OpenStudy (wolf1728):

I guess that is a good summary. I just don't like that expression of "this results in nothing". (Doesn't sound very mathematical).

OpenStudy (anonymous):

Loll any other way of putting it?

OpenStudy (wolf1728):

I suppose not.

OpenStudy (anonymous):

How about it sums up to nothing or zero?

OpenStudy (wolf1728):

Well I guess you could say that

OpenStudy (anonymous):

So you think this is all I need to as the answer?

OpenStudy (wolf1728):

Well looking at the question at the top: *** solution to the system It has no solution ***the two equations they are equivalent ***the result if the system was solved algebraically It adds to nothing - there is no algebraic solution.

OpenStudy (anonymous):

So should I write this instead and take out the other ones?

OpenStudy (wolf1728):

I'll leave that up to you - it's your class, your grade, etc

OpenStudy (anonymous):

Well you kind of already answered it for me I'm just saying should i add the examples you were giving me

OpenStudy (wolf1728):

At least the equivalent equation example. The inconsistent example is really not covered in that question.

OpenStudy (anonymous):

Thats cool. this is what i put now f the graph of both lines lie one on top of the other, there will be no solution.The equations are going to be equivalent. If solved algebraically the result will be nothing. when solving equivalent equations: A) x + y = 9 B) 3x + 3y = 27 We'll multiply equation A by -3 A) -3x -3y = -27 and add that to B B) 3x + 3y = 27 which sums up to nothing.

OpenStudy (wolf1728):

No the equations ARE equivalent not they are going to be ... Change "when solving equivalent equations" to "Example of trying to solve equivalent equations"

OpenStudy (anonymous):

oh ok

OpenStudy (wolf1728):

all set then?

OpenStudy (anonymous):

Yeah man Thanks big time

OpenStudy (anonymous):

I can always count on you

OpenStudy (wolf1728):

that's okay :-)

OpenStudy (anonymous):

:) Goodnight

OpenStudy (wolf1728):

good night